EA010A

PI: Alef
Title: AD Leo

Session: Oct 1996 (session 3)
Frequency: 3.6 cm
Mode: A
Obs time (UT): 303 EU 02:10(29/10)-12:00(29/10)
Stations: EbNtOnShUrWb
Processor: Bonn
EU Mk3 Tapes/St.: 2

Comment


To give feedback on this experiment, click here

Experiment Feedback

HartRAO (Jonathan Quick at Tue Oct 29 06:19:32 EST 1996) Minor failures
Tape splice caused vacuum loop dropout on obs at 08:22 UT

Onsala (J. Conway, A. Polatidis at Wed Oct 30 03:58:38 EST 1996) Success
No known problems

Shanghai (Huang Xinyong at Tue Nov 5 03:51:10 EST 1996) Success
No problems. Clock offset 0.0 micro sec. Weather was cloudy.

OAN\-Yebes (Pablo de Vicente at Tue Nov 5 08:30:35 EST 1996) Minor failures
FS computer problems. Observations lost on day 309 for some periods: 02:05:00 - 02:36:36, 06:15:00 - 06:40:16 and 06:54:00 - 07:35:36

Noto (Corrado Trigilio at Tue Nov 5 11:25:59 EST 1996) Success
No problems reported. Good sky conditions.

Effelsberg (D Graham at Wed Nov 6 09:16:28 EST 1996) Success
no known problems

Urumqi (Liu Xiang at Tue Nov 19 03:42:45 EST 1996) Minor failures
UT02:05:00-02:50:30, AZ and EL have errors. UT07:33:00-07:55:36, MK3 computer dead. Clock offset 0.1+/-0.5us.

MPIfR (Walter Alef at Tue Sep 23 17:23:17 METDST 1997) Minor failures
First tape of Urumqi no fringes, but rest of tapes O.K.. Lot of missing data for Crimea due to no power: only 2 hours of data. About 1 hours lost in Yebes. Fringes for: Effelsberg, Shanghai, Crimea, Noto, Hartrao, Onsala, Wettzell, Yebes, Madrid.